OpenStudy (anonymous):

1/9 + 1/2x = 1/x^2 solve.

OpenStudy (jiteshmeghwal9):

\[\huge{\frac{1}{9}+\frac{1}{2x}={1 \over x^2}}\]\[\Huge{\frac{2x+9}{18x}=\frac{1}{x^2}}\]\[\Huge{x^2(2x+9)=18x}\]\[\Huge{2x^3+9x^2=18x}\]\[\Huge{2x^3+9x^2-18x=0}\]find the roots f the equation :)

OpenStudy (jiteshmeghwal9):

of*

OpenStudy (anonymous):

thankyou!